# PO-Flow: Flow-based Generative Models for Sampling Potential Outcomes and Counterfactuals

### v2

@@ Appendix D Additional Experimental Results
D.1 Empirical Validation of Assumption (A3)
@@
Table 7: Empirical validation of Assumption (A3), $p_{\theta}(Z\mid X,A)=q(Z)=\mathcal{N}(0,1)$. We report two-sample MMD values comparing the empirical joint distribution $\{(z_{i},x_{i},a_{i})\}$ with synthetic samples $\{(z_{i}^{\prime},x_{i},a_{i})\}$ across datasets. Lower MMD indicates closer agreement with the assumed factorization $p(Z,X,A)=q(Z)p(X,A)$.
@@
As shown in Table 7, PO-Flow achieves small two-sample joint MMD values, which are comparable to those obtained from the ground-truth independent samples $\{(z_{i}^{\prime},x_{i},a_{i})\}$ and $\{(z_{i}^{\prime\prime},x_{i},a_{i})\}$. This provides empirical support for the validity of Assumption (A3).
D.2 Identifying the Most Likely Potential Outcomes via Log-Density
As shown in Proposition 4.7 and Section 5.3.1, we can naturally select the most likely potential outcome based on log-density. Table 8 reports the RMSE of PO-Flow when using the most likely PO selected by the algorithm. Specifically, for each sample we simulate 100 POs with PO-Flow, select the one with the highest log-density, and compute the results accordingly.
Table 8: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets. For PO-Flow, results are based on the most likely PO selected from 100 simulated outcomes using log-density. All values are averaged over 10-fold cross-validation.
@@ D.2 Identifying the Most Likely Potential Outcomes via Log-Density
Compared with the averaged RMSE over all POs generated by PO-Flow in Table 1, selecting the most likely POs based on log-density yields smaller RMSEs, providing more confident and closer estimates.
D.3 Convergence and Computational Costs
@@
Table 9: Comparison of model size, training time untill convergence, and sampling cost (for 500 samples).
@@
In addition, Table 9 compares model size, training time to convergence, and sampling time (for 500 samples) across PO-Flow and other deep generative baselines. All experiments were conducted on a single A100 GPU. Notably, PO-Flow attains superior results with the fewest neural network parameters, leading to faster training and substantially reduced sampling time compared with diffusion-, discrete flow-, and GAN-based methods.

### v3

@@ Appendix D Additional Experimental Results
D-A Empirical Validation of Assumption (A3)
@@
TABLE VII: Empirical validation of Assumption (A3), $p_{\theta}(Z\mid X,A)=q(Z)=\mathcal{N}(0,1)$. We report two-sample MMD values comparing the empirical joint distribution $\{(z_{i},x_{i},a_{i})\}$ with synthetic samples $\{(z_{i}^{\prime},x_{i},a_{i})\}$ across datasets. Lower MMD indicates closer agreement with the assumed factorization $p(Z,X,A)=q(Z)p(X,A)$.
@@
As shown in Table VII, PO-Flow achieves small two-sample joint MMD values, which are comparable to those obtained from the ground-truth independent samples $\{(z_{i}^{\prime},x_{i},a_{i})\}$ and $\{(z_{i}^{\prime\prime},x_{i},a_{i})\}$. This provides empirical support for the validity of Assumption (A3).
D-B Identifying the Most Likely Potential Outcomes via Log-Density
As shown in Proposition V.7, we can naturally select the most likely potential outcome based on log-density. Table VIII reports the RMSE of PO-Flow when using the most likely PO selected by the algorithm. Specifically, for each sample we simulate 100 POs with PO-Flow, select the one with the highest log-density, and compute the results accordingly.
TABLE VIII: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets. For PO-Flow, results are based on the most likely PO selected from 100 simulated outcomes using log-density. All values are averaged over 10-fold cross-validation.
@@ D-B Identifying the Most Likely Potential Outcomes via Log-Density
Compared with the averaged RMSE over all POs generated by PO-Flow in Table I, selecting the most likely POs based on log-density yields smaller RMSEs, providing more confident and closer estimates.
D-C Convergence and Computational Costs
@@
TABLE IX: Comparison of model size, training time until convergence, and sampling cost (for 500 samples).
@@
In addition, Table IX compares model size, training time to convergence, and sampling time (for 500 samples) across PO-Flow and other deep generative baselines. All experiments were conducted on a single A100 GPU. Notably, PO-Flow attains superior results with the fewest neural network parameters, leading to faster training and substantially reduced sampling time compared with diffusion-, discrete flow-, and GAN-based methods.